# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 5.1. Main estimates
(5.8) $\begin{split}\|X_{t}^{(n)}\|^{\beta+2}_{H}e^{\gamma t}=&\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma s}+\frac{\beta+2}{2}\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\left(2\langle A(X_{r}^{(n)}),X_{r}^{(n)}\rangle\right)\,\mathrm{d}r\\ &+\frac{\beta+2}{2}\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\|Q_{n}BQ_{n}\|_{2}^{2}\,\mathrm{d}r\\ &+\beta(\beta+2)\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta-2}_{H}\|(Q_{n}BQ_{n})^{\ast}X^{(n),x_{n}}_{r}\|^{2}_{H}\,\mathrm{d}r\\ &+(\beta+2)\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle\\ &+(\beta+2)\int_{s}^{t}\int_{D^{c}}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X^{(n),x_{n}}_{r},Q_{n}G(w)\rangle\,\widetilde{N}(\mathrm{d}r,\mathrm{d}w)\\ &+\int_{s}^{t}\int_{D^{c}}e^{\gamma r}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|^{\beta+2}_{H}-\|X_{r}^{(n)}\|^{\beta+2}_{H}\right.\\ &\qquad\left.-(\beta+2)\|X_{r}^{(n)}\|_{H}^{\beta}\langle X_{r}^{(n)},Q_{n}G(w)\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)\\ &+\gamma\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r\end{split}$
for all $0\leq s\leq t\leq T$, where $\widetilde{N}$ denotes the compensated Poisson random measure defined in (2.3). And thus, for any $0\leq s\leq t\leq T$, $\mathbb{P}$-a.s., by Hypothesis (B), noting that $B$ is symmetric,
$\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}+\delta_{1}\frac{\beta+2}{2}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}$
$+\frac{\beta+2}{2}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r$
$+(\beta+2)e^{-\gamma(t\wedge\tau_{R})}\sup_{s\leq u\leq t\wedge\tau_{R}}\left|\int_{s}^{u}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle\right|$
$+(\beta+2)e^{-\gamma(t\wedge\tau_{R})}\sup_{s\leq u\leq t\wedge\tau_{R}}\left|\int_{s}^{u}\int_{D^{c}}e^{\gamma r}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}G(w)\rangle\,\widetilde{N}(\mathrm{d}r,\mathrm{d}w)\right|$
$+\beta(\beta+2)\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(H)}\,\mathrm{d}r$
$+\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r,$
where the penultimate term is non-negative by the subgradient inequality (see e.g. [24]) and the fact that $u\mapsto\|u\|^{\beta+2}_{H}$ is convex with Fréchet differential $u\mapsto(\beta+2)\|u\|^{\beta}_{H}\langle u,h\rangle$, $h\in H$. By the Burkholder-Davis-Gundy inequality [45, Theorem 1], and by Young’s inequality,
$\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]+\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r$
$\leq$ $\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}\right]$
$+\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r$
$+3(\beta+2)\mathbb{E}\left[\left(\int_{s}^{t\wedge\tau_{R}}e^{2\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|B\|_{2}^{2}\,\mathrm{d}r\right)^{\frac{1}{2}}\right]$
$+3(\beta+2)\mathbb{E}\left[\left(\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{2\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|G(w)\|_{H}^{2}\,\pi(\mathrm{d}w)\,\mathrm{d}r\right)^{\frac{1}{2}}\right]$
$+\beta(\beta+2)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(H)}\,\mathrm{d}r$
$+\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r$
$\leq$ $\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}\right]$
$+\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r$
$+9(\beta+2)^{2}\mathbb{E}\left[\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|_{2}^{2}\,\mathrm{d}r\right]+\frac{1}{4}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]$
$+9(\beta+2)^{2}\mathbb{E}\left[\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|G(w)\|_{H}^{2}\,\pi(\mathrm{d}w)\,\mathrm{d}r\right]$
$+\frac{1}{4}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]$
$+\beta(\beta+2)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(H)}\,\mathrm{d}r$
$+\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r,$
Hence, for $\lambda_{0}\in(0,1)$ as in Definition 5.1, and by Fubini’s theorem,
$\frac{1}{2}\mathbb{E}\left[\sup_{s\leq r\leq t\wedge\tau_{R}}\|X_{r}^{(n)}\|^{\beta+2}_{H}\right]$
$+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r+(1-\lambda_{0})I_{t\wedge\tau_{R},s}^{(n)}$
$\leq$ $\mathbb{E}\left[\|X_{s}^{(n)}\|_{H}^{\beta+2}e^{\gamma(s-(t\wedge\tau_{R}))}\right]$
$+\left(\left(\frac{1}{2}(\beta+2)+9(\beta+2)^{2}\right)\|B\|_{2}^{2}+\beta(\beta+2)\|B\|_{L(H)}^{2}\right)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r$
$+9(\beta+2)^{2}\mathfrak{g}_{0}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r$
$+\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.$
$\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,N(\mathrm{d}r,\mathrm{d}w)$
$+\gamma\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r,$
where
$I_{t,s}^{(n)}:$ $=c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}e^{-\gamma t}\mathbb{E}\int_{s}^{t}e^{\gamma r}\|X_{r}^{(n)}\|^{\alpha+\beta}_{H}\,\mathrm{d}r,$
for any $0\leq s\leq t\leq T$ and any $n\in\mathbb{N}$.
@@
Our hypotheses guarantee that all constants in Definition 5.1 are finite. Recall Young’s inequality for products, that is, for any $p,q\in(1,\infty)$ satisfying $p^{-1}+q^{-1}=1$ it follows that
@@
If $\alpha>2$, we get that for any $x\in H$, and any $\lambda_{i}\in[0,1)$, $i=1,2,3,4$, and $\lambda_{0}\in(0,1]$ with $\sum_{i=0}^{4}\lambda_{i}=1$, $c_{i}\geq 0$, $i=1,\ldots,5$ as in Definition 5.1, with
@@
(5.10) $\begin{split}&\left(\left(\frac{\beta+2}{2}+9(\beta+2)^{2}\right)\|B\|_{2}^{2}+\beta(\beta+2)\|B\|_{L(H)}^{2}\right)\|x\|^{\beta}_{H}\leq\lambda_{1}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{1},\\ &\gamma\|x\|^{\beta+2}_{H}\|B\|_{2}^{2}\leq\lambda_{2}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{2},\\ &9(\beta+2)^{2}\mathfrak{g}_{0}\|x\|^{\beta}_{H}\leq\lambda_{1}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{3},\\ &C_{\beta}\mathfrak{g}_{0}\|x\|^{\beta}_{H}\leq\lambda_{4}c_{0}^{\alpha}\delta_{1}\frac{\beta+2}{2}\|x\|^{\alpha+\beta}_{H}+c_{4},\\ &C_{\beta}\mathfrak{g}_{\beta}=c_{5},\end{split}$
for any $\gamma>0$. If $\alpha=2$, the inequalities for $c_{1},c_{3}$, and $c_{4}$ follow from (5.9), and $c_{2}=0$ follows by choosing $\gamma\in(0,\delta_{2})$ such that
(5.11) $\gamma\|B\|_{2}^{2}\leq\lambda_{2}c_{0}^{2}\delta_{1}\frac{\beta+2}{2}.$
As a consequence, we have that for $c_{i}\geq 0$, $i=1,2,3$, as in Definition 5.1, that $\mathbb{P}$-a.s.
(5.12) $\begin{split}&\left(\left(\frac{\beta+2}{2}+9(\beta+2)^{2}\right)\|B\|_{2}^{2}+\beta(\beta+2)\|B\|_{L(H)}^{2}\right)\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r\\ \leq&\lambda_{1}I_{{t\wedge\tau_{R}},s}^{(n)}+c_{1}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r,\\ &\gamma\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta+2}_{H}\,\mathrm{d}r\leq\lambda_{2}I_{{t\wedge\tau_{R}},s}^{(n)}+c_{2}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r,\\ &9(\beta+2)^{2}\mathfrak{g}_{0}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|^{\beta}_{H}\,\mathrm{d}r\leq\lambda_{3}I_{{t\wedge\tau_{R}},s}^{(n)}+c_{3}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r,\end{split}$
We claim that $\mathbb{P}$-a.s.
(5.13) $\begin{split}&\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.\\ &\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,\pi(\mathrm{d}w)\,\mathrm{d}r\\ \leq&\lambda_{4}I_{{t\wedge\tau_{R}},s}^{(n)}+(c_{4}+c_{5})\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r.\end{split}$
By Lemma A.1, we have for $C_{\beta}=((2^{\beta-3}\vee 1)\beta(\beta+2))\vee((2^{\beta-1}\vee 1)(\beta+2))$, $\beta>0$, and $C_{0}=\frac{1}{2}$ that
$\|x+w\|_{H}^{\beta+2}-\|x\|^{\beta+2}_{H}-(\beta+2)\|x\|^{\beta}_{H}\left\langle x,w\right\rangle\leq C_{\beta}\left(\|x\|_{H}^{\beta}\|w\|_{H}^{2}+\|w\|^{\beta+2}_{H}\right).$
Hence, by Fubini’s theorem,
(5.14) $\begin{split}&\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}+Q_{n}G(w)\|_{H}^{\beta+2}\right.\\ &\qquad\left.-\|X_{r}^{(n)}\|^{\beta+2}_{H}-(\beta+2)\|X_{r}^{(n)}\|^{\beta}_{H}\left\langle X_{r}^{(n)},Q_{n}G(w)\right\rangle\right]\,\pi(\mathrm{d}w)\,\mathrm{d}r\\ \leq&C_{\beta}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}\int_{D^{c}}e^{\gamma(r-(t\wedge\tau_{R}))}\left[\|X_{r}^{(n)}\|_{H}^{\beta}\|Q_{n}G(w)\|_{H}^{2}+\|Q_{n}G(w)\|^{\beta+2}_{H}\right]\,\pi(\mathrm{d}w)\,\mathrm{d}r\\ \leq&C_{\beta}\mathfrak{g}_{0}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\|X_{r}^{(n)}\|_{H}^{\beta}\,\mathrm{d}r+C_{\beta}\mathfrak{g}_{\beta}\mathbb{E}\int_{s}^{t\wedge\tau_{R}}e^{\gamma(r-(t\wedge\tau_{R}))}\,\mathrm{d}r.\end{split}$
The claim follows now by (5.10). Thus, noting that, by definition, $1-\lambda_{0}=\sum_{i=1}^{4}\lambda_{i}$, and recalling that $\mathfrak{c}:=\sum_{i=1}^{5}c_{i}$,

### v4

@@ 5.1. Main estimate
(5.8) $\begin{split}\|X_{t}^{(n)}\|^{\beta+2}_{H}=&\|X_{s}^{(n)}\|_{H}^{\beta+2}+\frac{\beta+2}{2}\int_{s}^{t}2\|X_{r}^{(n)}\|^{\beta}_{H}\langle A(X_{r}^{(n)}),X_{r}^{(n)}\rangle\,\mathrm{d}r\\ &+\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|Q_{n}B\tilde{Q}_{n}\|^{2}_{L_{2}(U,H)}\,\mathrm{d}r\\ &+\beta(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta-2}_{H}\|(Q_{n}B\tilde{Q}_{n})^{\ast}X^{(n)}_{r}\|^{2}_{U}\,\mathrm{d}r\\ &+(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle,\end{split}$
for all $0\leq s\leq t\leq T$. And thus, for any $0\leq s\leq t\leq T$, $\mathbb{P}$-a.s., by Hypothesis (B), and Fact 3.1, for $\lambda_{0}\in(0,1)$,
$\|X_{t}^{(n)}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r+(1-\lambda_{0})\delta_{1}c_{0}^{\alpha}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\alpha+\beta}_{H}\,\mathrm{d}r$
$\leq$ $\|X_{t}^{(n)}\|^{\beta+2}_{H}+\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{(n)}\|_{H}^{\beta+2}+\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|Q_{n}B\tilde{Q}_{n}\|^{2}_{L_{2}(U,H)}\,\mathrm{d}r$
$+\beta(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta-2}_{H}\|(Q_{n}B\tilde{Q}_{n})^{\ast}X^{(n)}_{r}\|^{2}_{U}\,\mathrm{d}r$
$+(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle$
$=$ $\|X_{s}^{(n)}\|_{H}^{\beta+2}+\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|Q_{n}B\tilde{Q}_{n}\|^{2}_{L_{2}(U,H)}\,\mathrm{d}r$
$+\beta(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta-2}_{H}\|(Q_{n}B\tilde{Q}_{n})^{\ast}X^{(n)}_{r}\|^{2}_{U}\,\mathrm{d}r$
$+(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle-\frac{(\beta+2)^{2}}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|Q_{n}B\tilde{Q}_{n}\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r$
$+\frac{(\beta+2)^{2}}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|Q_{n}B\tilde{Q}_{n}\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{(n)}\|_{H}^{\beta+2}+\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L_{2}(U,H)}\,\mathrm{d}r$
$+\beta(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|B\|^{2}_{L(U,H)}\,\mathrm{d}r$
$+(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle-\frac{(\beta+2)^{2}}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|Q_{n}B\tilde{Q}_{n}\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r$
$+\frac{(\beta+2)^{2}}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|B\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r.$
@@
Recall Young’s inequality for products, that is, for any $p,q\in(1,\infty)$ satisfying $p^{-1}+q^{-1}=1$ it follows that
@@
For any $y\in H$, and any $\lambda_{i}\in(0,1)$, $i=1,2,3$, and $\lambda_{0}\in(0,1)$ with $\sum_{i=0}^{3}\lambda_{i}=1$, and $c_{i}\geq 0$, $i=1,2,3$ as follows, compare also with Hypothesis (H),
@@
(5.10) $\begin{split}&c_{1}:=\frac{\alpha(\beta+2)}{2(\alpha+\beta)\left(\lambda_{1}\delta_{1}c_{0}^{\alpha}\frac{\alpha+\beta}{\beta}\right)^{\frac{\beta}{\alpha}}}\|B\|_{L_{2}(U,H)}^{\frac{2(\alpha+\beta)}{\alpha}},\\ &c_{2}:=\beta^{\frac{\alpha+\beta}{\alpha}}\frac{\alpha(\beta+2)}{(\alpha+\beta)\left(\lambda_{2}\delta_{1}c_{0}^{\alpha}\frac{\alpha+\beta}{2\beta}\right)^{\frac{\beta}{\alpha}}}\|B\|_{L(U,H)}^{\frac{2(\alpha+\beta)}{\alpha}},,\\ &c_{3}:=\frac{(\beta+2)(\alpha-\beta-2)}{2(\alpha+\beta)\left(\lambda_{3}\delta_{1}c_{0}^{\alpha}\frac{\alpha+\beta}{2\beta+2}\right)^{\frac{2\beta+2}{\alpha-\beta-2}}}(\beta+2)^{\frac{\alpha+\beta}{\alpha-\beta-2}}\|B\|_{L_{2}(U,H)}^{\frac{2(\alpha+\beta)}{\alpha-\beta-2}},\end{split}$
we get the following inequalities by (5.9), if $\beta>0$,
(5.11) $\begin{split}&\frac{\beta+2}{2}\|y\|_{H}^{\beta}\|B\|_{L_{2}(U,H)}^{2}\leq\lambda_{1}\frac{\beta+2}{2}\delta_{1}c_{0}^{\alpha}\|y\|_{H}^{\alpha+\beta}+c_{1},\\ &(\beta+2)\beta\|y\|^{\beta}_{H}\|B\|^{2}_{L(U,H)}\leq\lambda_{2}\frac{\beta+2}{2}\delta_{1}c_{0}^{\alpha}\|y\|_{H}^{\alpha+\beta}+c_{2},\end{split}$
and otherwise, if $\beta=0$, then $\lambda_{1}=\lambda_{2}=0$ and,
(5.12) $\begin{split}&c_{1}=\|B\|_{L_{2}(U,H)}^{2}\\ &c_{2}=0,\end{split}$
furthermore, if $\alpha+\beta>2\beta+2$,
(5.13) $\begin{split}&\frac{(\beta+2)^{2}}{2}\|y\|_{H}^{2\beta+2}\|B\|_{L_{2}(U,H)}^{2}\leq\lambda_{3}\frac{\beta+2}{2}\delta_{1}c_{0}^{\alpha}\|y\|_{H}^{\alpha+\beta}+c_{3},\end{split}$
and otherwise, if $\alpha=\beta+2$, then $c_{3}=0$, and by assumption,
(5.14) $\begin{split}&\|B\|_{L_{2}(U,H)}^{2}\leq\lambda_{3}\frac{1}{\alpha}\delta_{1}c_{0}^{\alpha},\end{split}$
which, in this case, is equivalent to
(5.15) $\begin{split}&\frac{(\beta+2)^{2}}{2}\|B\|_{L_{2}(U,H)}^{2}\leq\lambda_{3}\frac{\beta+2}{2}\delta_{1}c_{0}^{\alpha}.\end{split}$
Combining these inequalities, we have that $\mathbb{P}$-a.s.
$\|X_{t}^{(n)}\|^{\beta+2}_{H}+\lambda_{0}\delta_{1}\frac{\beta+2}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\|X_{r}^{(n)}\|^{\alpha}_{V}\,\mathrm{d}r$
$\leq$ $\|X_{s}^{(n)}\|_{H}^{\beta+2}+(t-s)(c_{1}+c_{2}+c_{3})$
$+(\beta+2)\int_{s}^{t}\|X_{r}^{(n)}\|^{\beta}_{H}\langle X_{r}^{(n)},Q_{n}B\,\mathrm{d}W_{r}^{(n)}\rangle-\frac{(\beta+2)^{2}}{2}\int_{s}^{t}\|X_{r}^{(n)}\|^{2\beta+2}_{H}\|Q_{n}B\tilde{Q}_{n}\|_{L_{2}(U,H)}^{2}\,\mathrm{d}r.$
Stopping and taking the exponential yields $\mathbb{P}$-a.s.,